A wire has a diameter of 5mm, length 50m and resistivity 3.2x10 to the power of -8 ohms meter. What is the resistance of my wire?
1
Expert's answer
2019-05-13T10:27:37-0400

The electrical resistance of a homogeneous conductor with a specific resistance ρ, a length l and a cross-sectional area S can be calculated by the formula

"R=\rho*l/S"

"S=\pi*d^2/4"

"R=4*\rho*l/(\pi*d^2)=0.082Ohm"

Answer: 0.082Ohm